# Video-MME: The First-Ever Comprehensive Evaluation Benchmark of Multi-modal LLMs in Video Analysis

### v1

@@ 3.3 Analysis
Most of evaluations only take video frames as input, requiring models to answer questions solely based on visual contexts. However, many videos inherently contain extra information from other modalities, such as subtitles and audios. To understand their impact, we vary the combinations of input modalities in the evaluation, and report results in Table 5 and Figure 4. We can draw the following observations. (1) Introducing subtitles and audios can improve the results. For example, in the multilingual task shown in Table 5, with the addition of audios, Gemini 1.5 Pro achieves +16.6% accuracy on long videos compared to the frame-only setting. This indicates that subtitles and audios provide some necessary information to answer the questions. (2) Subtitles and audios provide greater assistance in understanding long videos compared to short videos. For example, in Table 5, compared to only using frames, the addition of subtitles improves the model’s performance by 2.4% on short videos and by 8.8% on long videos. This is because test samples of long videos include more challenging reasoning questions, which requires the model to utilize subtitles and audio information for accurate responses. (3) For MLLMs, using subtitles is more effective than audio. Subtitles are usually transcriptions of audio, primarily capturing speech content, while audios encompass more ambient sounds. As can be seen from Table 5, whether it is short, medium, or long videos, subtitles bring higher improvement than audios. There are some exceptions in Table 5 and Figure 4, such as Multilingual, which may be due to the quality of the subtitles themselves. In addition, the audio also includes some content that subtitles can not express, such as singing and intonation.
@@
In Table 4, we respectively compare the performance of different models on short, medium, and long videos. As video duration increases, both open-sourced and commercial models exhibit a significant decline in performance. For example, the accuracy of LLaVA-NeXT-Video drops by 12% and 18.5% from short to medium and long videos, respectively, while Gemini 1.5 Pro’s accuracy decreases by 7% and 14.8%. There are three main reasons for the performance decline. (1) Increased proportion of difficult tasks. As shown in the bottom right corner of Figure 2, test samples for long videos contain a higher proportion of reasoning questions. These questions are more challenging than perception and recognition tasks, thus posing a greater challenge to the model’s capabilities. (2) Increased sparsity in frame sampling, leading to a reduction in effective input information. Ideally, for videos of varying lengths, models should sample video frames at a fixed fps to ensure consistent information density in frame sequences [53, 52]. However, many open-sourced models [23, 28] fix the number of input frames, e.g., 8 frames, resulting in excessively sparse information density as the video length increases. This sparsity prevents the model from retaining all useful visual semantics, hindering accurate predictions. Introducing additional modalities, e.g., subtitles, can effectively supplement the missing information [7]. (3) Increased difficulty in long context understanding. Although Gemini 1.5 Pro correspondingly increases the number of sampled frames in the long video, there is still a significant performance degradation. Understanding the long context of either single-modality (LLM) or multi-modality (MLLM) is always a great challenge.
@@ 5 Related Work
Alongside advancements in architecture, significant efforts have been made to improve benchmarking for MLLMs, guiding the development of the next generation of these models. Previous studies have integrated various aspects of evaluation, such as perception and cognitive capabilities, to create comprehensive benchmarks for assessing image MLLMs [13, 61, 35]. As image MLLMs have demonstrated exceptional performance in general perception tasks, benchmarks regarding scientific understanding [26], multi-modal mathematical reasoning [38, 67], and multi-disciplinary [63] capabilities have drawn increasing attention. For video MLLMs, similar efforts have been made to incorporate existing benchmarks for evaluating video understanding [24, 44]. Given the temporal nature of video modalities, specific benchmarks have been developed to address temporal understanding, highlighting the limitations of current video MLLMs in comprehending video content [27, 36].

### v2

@@ 3.3 Analysis
Most of evaluations only take video frames as input, requiring models to answer questions solely based on visual contexts. However, many videos inherently contain extra information from other modalities, such as subtitles and audios. To understand their impact, we vary the combinations of input modalities in the evaluation, and report results in Table 5 and Figure 4. We can draw the following observations. (1) Introducing subtitles and audios can improve the results. For example, in the multilingual task shown in Table 5, with the addition of subtitles/audios, Gemini 1.5 Pro achieves +16.7%/+12.5% accuracy on long videos compared to the frame-only setting. This indicates that subtitles and audios provide some necessary information to answer the questions. (2) Subtitles and audios provide greater assistance in understanding long videos compared to short videos. For example, in Table 5, compared to only using frames, the addition of subtitles improves the model’s performance by 2.8% on short videos and by 10.1% on long videos. This is because test samples of long videos include more challenging reasoning questions, which requires the model to utilize subtitles and audio information for accurate responses. (3) For MLLMs, using subtitles is more effective than audios. Subtitles are usually transcriptions of audio, primarily capturing speech content, while audios encompass more ambient sounds. As can be seen from Table 5, whether it is short, medium, or long videos, subtitles bring higher improvement than audios. There are some exceptions in Table 5 and Figure 4, such as Multilingual, which may be due to the quality of the subtitles themselves. In addition, the audio also includes some content that subtitles can not express, such as singing and intonation.
@@
In Table 4, we respectively compare the performance of different models on short, medium, and long videos. As video duration increases, both open-sourced and commercial models exhibit a significant decline in performance. For example, the accuracy of VILA-1.5 drops by 10% and 17.3% from short to medium and long videos, respectively, while Gemini 1.5 Pro’s accuracy decreases by 7.4% and 14.3%. There are three main reasons for the performance decline. (1) Increased proportion of difficult tasks. As shown in the bottom right corner of Figure 2, test samples for long videos contain a higher proportion of reasoning questions. These questions are more challenging than perception and recognition tasks, thus posing a greater challenge to the model’s capabilities. (2) Increased sparsity in frame sampling, leading to a reduction in effective input information. Ideally, for videos of varying lengths, models should sample video frames at a fixed fps to ensure consistent information density in frame sequences [56, 55]. However, many open-sourced models fix the number of input frames, e.g., 8 frames, resulting in excessively sparse information density as the video length increases. This sparsity prevents the model from retaining all useful visual semantics, hindering accurate predictions. Introducing additional modalities, e.g., subtitles, can effectively supplement the missing information [7]. (3) Increased difficulty in long context understanding. Although Gemini 1.5 Pro correspondingly increases the number of sampled frames in the long video, there is still a significant performance degradation. Understanding the long context of either single-modality (LLM) or multi-modality (MLLM) is always a great challenge.
@@ 5 Related Work
Alongside advancements in architecture, significant efforts have been made to improve benchmarking for MLLMs, guiding the development of the next generation of these models. Previous studies have integrated various aspects of evaluation, such as perception and cognitive capabilities, to create comprehensive benchmarks for assessing image MLLMs [13, 67, 38]. As image MLLMs have demonstrated exceptional performance in general perception tasks, benchmarks regarding scientific understanding [28], multi-modal mathematical reasoning [41, 73], and multi-disciplinary [69] capabilities have drawn increasing attention. For video MLLMs, similar efforts have been made to incorporate existing benchmarks [59, 26] for evaluating video understanding [24, 47]. Given the temporal nature of video modalities, specific benchmarks have been developed to address temporal understanding, highlighting the limitations of current video MLLMs in comprehending video content [29, 39].